# Multilingual != Multicultural: Evaluating Gaps Between Multilingual Capabilities and Cultural Alignment in LLMs

### v1

@@ Abstract
Large Language Models (LLMs) are becoming increasingly capable across global languages. However, the ability to communicate across languages does not necessarily translate to appropriate cultural representations. A key concern is US-centric bias, where LLMs reflect US rather than local cultural values. We propose a novel methodology that compares LLM-generated response distributions against population-level opinion data from the World Value Survey across four languages (Danish, Dutch, English, and Portuguese). Using a rigorous linear mixed-effects regression framework, we compare two families of models: Google’s Gemma models (2B–27B parameters) and successive iterations of OpenAI’s turbo-series. Across the families of models, we find no consistent relationships between language capabilities and cultural alignment. While the Gemma models have a positive correlation between language capability and cultural alignment across languages, the OpenAI models do not. Importantly, we find that self-consistency is a stronger predictor of multicultural alignment than multilingual capabilities. Our results demonstrate that achieving meaningful cultural alignment requires dedicated effort beyond improving general language capabilities.
Multilingual != Multicultural: Evaluating Gaps Between Multilingual Capabilities and Cultural Alignment in LLMs
Jonathan Rystrøm  and Hannah Rose Kirk  and Scott Hale Oxford Internet Institute, University of Oxford, UK jonathan.rystrom@oii.ox.ac.uk
@@ 1 Introduction
[…]
Spearheaded by accessible chat interfaces to powerful models like ChatGPT (OpenAI, 2022), LLMs are reaching hundreds of millions of users (Milmo, 2023). These models are deployed across diverse contexts: from tutoring mathematics (Khan, 2023) to building software applications (Peng et al., 2023) to assisting in legal cases (Tan et al., 2023). While most LLMs demonstrate multilingual abilities (Üstün et al., 2024), the ability to communicate across languages does not necessarily translate into appropriate cultural representations. Disentangling language capabilities and cultural alignment is crucial for understanding how LLMs should be examined and audited (Mökander et al., 2024) and for ensuring these technologies work for diverse people (D’Ignazio and Klein, 2020; Weidinger et al., 2022).
A priori, we might expect these models to exhibit US-centric cultural biases despite their multilingual capabilities, given their development context: Although millions use LLMs, they are developed by a select few. Many model providers, such as OpenAI, Anthropic, and Google, are American technology companies headquartered in Silicon Valley. These companies comprise a narrow slice of human experience, limiting the voices that contribute to critical design decisions in LLMs (D’Ignazio and Klein, 2020). The companies typically train LLMs on massive amounts of predominantly English text and employ American crowd workers to rate and evaluate the LLMs’ responses (Johnson et al., 2022; Kirk et al., 2023). Far too often, the benefits and harms of data technologies are unequally distributed, reinforcing biases and harming already minoritized groups (Birhane, 2020; Milan and Treré, 2019; Khandelwal et al., 2024). Understanding how LLMs represent different cultures is thus paramount to establishing risks of representational harm (Rauh et al., 2022) and ensuring the technology’s utility is shared across diverse communities.
@@
Although studies like the World Values Survey (WVS) have documented how values vary across cultures (EVS/WVS, 2022), it remains unclear whether more capable LLMs—through scaling or improved training—better align with these cultural differences (Bai et al., 2022; Kirk et al., 2023). While the WVS has been used in prior research on values in LLMs, these studies have focused predominately on individual models’ performance within an English-language context. (Cao et al., 2023; Arora et al., 2023; AlKhamissi et al., 2024). This paper addresses this gap by developing a methodology for assessing how well families of LLMs represent different cultural contexts across multiple languages. We compare two distinct paths to model improvement: systematic scaling of instruction-tuned models and commercial product development comprising scaling and innovation in post-training to accommodate pressures from capabilities, cost, and preferences (OpenAI et al., 2024b).
@@
We operationalise multilingual capability as an LLM’s performance on a range of multilingual benchmarks across languages (see, e.g., Nielsen, 2023). We describe the specific benchmarks and performances in Appendix F.
This work makes several key contributions. First, we introduce a novel distribution-based methodology for probing cultural alignment across languages, moving beyond direct survey approaches to better capture underlying cultural values (Sorensen et al., 2024). Second, we provide the first systematic comparison of how improvements in scale and post-training affect cultural alignment and US-centric bias across English, Danish, Dutch, and Portuguese through a series of robust statistical models. Third, we release a dataset of model-generated responses across multiple languages and cultural contexts as well as our code, enabling future research into cultural alignment and bias. [^1] See https://github.com/jhrystrom/multicultural-alignment for code and https://huggingface.co/datasets/ryzzlestrizzle/multicultural-wvs-alignment for data Together, these contributions advance our understanding of how LLM development choices influence cultural representation while providing tools for ongoing investigation of these critical issues.
@@ 2 Measuring Cultural Alignment
Figure 2: Pearson correlations in value polarity scores across studied countries from the World Values Survey. All correlations are positive with most being between 0.7-0.95.
This section defines what we mean by ‘cultural alignment’ and how to measure it in LLMs. First, we provide a high-level intuition of the cultural alignment as reproducing distributions of values in a particular population. Then we show how to a) get a ground-truth distribution of values using the World Values Survey (§2.1) and b) elicit value distributions from LLMs (§2.2)
@@
Within a culture there will be a variety of stances to any particular topic. However, the distribution of stances will be characteristic among cultures. For instance, while some Danes are opposed to abortion, it is a much less contentious topic than in the US Adamczyk et al. (2020); Norup (1997).
We posit that cultural alignment for a specific group of people can be operationalised as how well an LLM reproduces the distribution of values over a wide range of topics (Sorensen et al., 2024). By investigating distributions of responses, we differ from previous work that directly surveys the LLMs as a regular participant (e.g., Cao et al., 2023). Our goal is to get more naturalistic elicitations of the underlying values and avoid sycophancy and response bias Sharma et al. (2023).
More specifically, we operationalise reproduction as having a high correlation between value polarity score: the fraction of people in favour of a topic in the population and the fraction of responses in favour of the topic elicited by the LLM. Note, that we binarise issues to allow for simpler operationalisation. Below, we describe how we empirically estimate the value polarity score for ground truth (§2.1) and LLMs (§2.2).
@@ 2.1 Ground Truth: World Values Survey
To get a ‘ground truth’ distribution of cultural values, we use the joint World Values Survey and European Values Survey (EVS; EVS/WVS, 2022). These large-scale international surveys cover adults across 92 countries with samples that are nationally representative for gender, age, education, and religion. The surveys’ broad coverage enables cross-cultural comparability, though some scholars note challenges in ensuring response comparability across countries (Alemán and Woods, 2016).
We select a subset of questions corresponding to values about the environment, work, family, politics & society, religion & morals, and security—all questions where there is are clear ‘for’ and ‘against’ positions. These questions comprise our topics. Appendix A shows the full list of included questions.
The final step is calculating the value polarity score. For each question, we define what an ‘affirmative’ response means (i.e., a Likert score higher/lower than the middle score, depending on the directionality of the question) and then define the value polarity score as the (weighted) fraction of respondents with an affirmative stance to the topic within the population. Thus, a culture’s values can be represented as a vector, where each element corresponds to a value polarity score for a specific topic.
@@ 2.2 Ecologically valid LLM responses
Developing ecologically valid prompts requires careful consideration. When evaluating LLM responses to value-laden topics, simply asking questions like “What proportion of people support topic X?” or “Do you support topic X?” proves inadequate (e.g., Hartmann et al., 2023). Such direct approaches suffer from three key limitations: they generate false positives through excessive agreement, fail to reflect realistic usage patterns, and provide insufficient variation to assess cultural alignment (Röttger et al., 2024). They also struggle to capture instance-specific harms that emerge when systems misalign with users’ cultural contexts (Rauh et al., 2022).
Instead, we adopt an implicit approach by asking the model to generate responses from hypothetical respondents. For example, prompting “imagine surveying 10 random people on topic X. What are their responses?” This method reveals the model’s latent opinion distribution while avoiding the limitations of direct questioning. Details for prompt construction are provided in Appendix B.
@@
Having a method for eliciting distributions of values, the next step is to seed culture. One typical way of seeding a specific culture is to explicitly instruct the LLM either by mentioning a specific country (‘imagine surveying 10 random Americans’) or through describing specific personas (‘Imagine surveying a 85 year old Danish woman…’; AlKhamissi et al., 2024). The problem with these approaches is that they stray from actual uses of LLMs. Users are unlikely to explicitly mention their demographic information or nationality (Zheng et al., 2023a).
Instead, we use language as a proxy for culture. For instance, a prompt in Danish is assumed to come from a Dane etc. For languages spoken in multiple countries this approach is intentionally ambiguous. The ambiguity allows us to elicit the underlying ‘default’ alignment rather than the general ability to emulate cultures (Tao et al., 2024). To create prompts across languages, we use gpt-3.5-turbo to translate our original English prompts, which we manually verify.
@@
Finally, to transform the LLMs’ hypothetical survey responses into vectors of stances, we use an LLM-as-a-judge approach Zheng et al. (2023b). Specifically, we use gpt-3.5-turbo (OpenAI, 2022) to label each substatement as either ‘pro’, ‘con’, or ‘null’ given the context of the topic. We then calculate the proportion of ‘pro’ versus ‘con’ responses as the LLM’s value polarity score for the given statement. These scores can then be used to compare against the value polarity scores from the WVS. Specifically, we can calculate the Spearman rank correlation to obtain a measure of similarity between the LLMs responses and the value distributions of a given population.
@@ 3 Experimental Setup
To investigate whether improving the multilingual capabilities of LLMs improves cultural alignment, we set up an experiment using a carefully chosen set of models and languages. We examine two different kinds of model improvements: scaling and commercial product development. These cases provide complementary perspectives on the effects of multilingual capabilities on cultural alignment. Scaling is the most well-studied path to improving LLMs (Kaplan et al., 2020; Ganguli et al., 2022). Commercial product development, on the other hand, comprises both scale and innovation in post-training to accommodate different pressures from capabilities, cost, and preferences (Kirk et al., 2024a). For scaling, we use the instruction-tuned Gemma models Gemma et al. (2024), while for product development, we use OpenAI’s turbo-series models OpenAI (2022); OpenAI et al. (2024a, b). We provide details of these model families in §3.1. A breakdown of the computational cost can be seen in Appendix E.
@@
For the languages, we compare English with Danish, Dutch, and Portuguese. This set allows us to test multiple assumptions about cultural alignment. English represents a widely-used case: it’s a global language with speakers across many countries represented in the WVS/EVS (see Fig. 2). This diversity allows us to assess whether LLMs align more strongly with US values or those of other English-speaking nations.
Danish and Dutch serve as controlled test cases since they are primarily used in single countries. If cultural alignment stems from pre-training data, models should show strong Danish/Dutch cultural alignment when using these languages, despite their small and low-quality share of training data (Kreutzer et al., 2022). Alternatively, if alignment emerges from English-based post-training processes (which are predominately English-based; Blevins and Zettlemoyer, 2022), responses in these languages should align more with US values.
@@
For each language/model combination we collect 300 prompt-response pairs to sufficiently power our statistical analysis (see §3.2). After filtering out responses that either lacked the required hypothetical survey format or were in another language than the prompt, we obtained between 111–299 valid responses per combination. We calculate the correlation in value polarity scores at three levels: country (e.g., US or Denmark), language (pooling all speakers of a given language), and global (weighted values from all WVS/EVS participants).

### v2

@@ Abstract
Large Language Models (LLMs) are becoming increasingly capable across global languages. However, the ability to communicate across languages does not necessarily translate to appropriate cultural representations. A key concern is US-centric bias, where LLMs reflect US rather than local cultural values. We propose a novel methodology that compares LLM-generated response distributions against population-level opinion data from the World Value Survey across four languages (Danish, Dutch, English, and Portuguese). Using a rigorous linear mixed-effects regression framework, we compare three families of models: Google’s Gemma models (2B–27B parameters), AI2’s OLMo models (7B-32B parameters), and successive iterations of OpenAI’s turbo-series. Across the families of models, we find no consistent relationships between language capabilities and cultural alignment. While the Gemma models have a positive correlation between language capability and cultural alignment across all languages, the OpenAI and OLMo models are inconsistent. Our results demonstrate that achieving meaningful cultural alignment requires dedicated effort beyond improving general language capabilities.
@@ 1 Introduction
Figure 1: The relationship between multilingual capability and cultural alignment is inconsistent across LLM families, as shown by coefficients from our linear mixed-effects model ($\beta_{multilingual}=\beta_{flm}$; Eq. 3; §3.2). OpenAI and OLMo models show negative or insignificant relationships outside of Danish and Dutch, while Gemma models show positive relationships throughout ($p<.05$).
Spearheaded by accessible chat interfaces to powerful models like ChatGPT (OpenAI, 2022), LLMs are reaching hundreds of millions of users (Milmo, 2023). These models are deployed across diverse contexts: from tutoring mathematics (Khan, 2023) to building software applications (Peng et al., 2023) to assisting in legal cases (Tan et al., 2023). While most LLMs demonstrate multilingual abilities (Üstün et al., 2024), the ability to communicate across languages does not necessarily translate into appropriate cultural representations. Disentangling language capabilities and cultural alignment is crucial for understanding how LLMs should be examined and audited (Mökander et al., 2024) and for ensuring these technologies work for diverse people (D’ignazio and Klein, 2023; Weidinger et al., 2022).
Given the Silicon Valley origins of many frontier AI labs and the prevalence of American English training data, we might expect LLMs to exhibit US-centric cultural biases despite their multilingual capabilities. These companies comprise a narrow slice of human experience, limiting the voices that contribute to critical design decisions in LLMs (D’ignazio and Klein, 2023). They typically train LLMs on massive amounts of predominantly English text and employ American crowd workers to rate and evaluate the LLMs’ responses (Johnson et al., 2022; Kirk et al., 2023). Far too often, the benefits and harms of data technologies are unequally distributed, reinforcing biases and harming already minoritized groups (Birhane, 2020; Milan and Treré, 2019; Khandelwal et al., 2024). Understanding how LLMs represent different cultures is thus paramount to establishing risks of representational harm (Rauh et al., 2022) and ensuring the technology’s utility is shared across diverse communities.
@@
Although studies like the World Values Survey (WVS) have documented how values vary across cultures (EVS/WVS, 2022), it remains unclear whether more capable LLMs—through scaling or improved training—better align with these cultural differences (Bai et al., 2022; Kirk et al., 2023). While the WVS has been used in prior research on values in LLMs, these studies have focused predominantly on individual models’ performance within an English-language context. (Cao et al., 2023; Arora et al., 2023; AlKhamissi et al., 2024). This paper addresses this gap by developing a methodology for assessing how well families of LLMs represent different cultural contexts across multiple languages. We compare two distinct paths to model improvement: systematic scaling of instruction-tuned models and commercial product development comprising scaling and innovation in post-training to accommodate pressures from capabilities, cost, and preferences (OpenAI et al., 2024b).
@@
We operationalise multilingual capability as an LLM’s performance on a range of multilingual benchmarks across languages (see, e.g., Nielsen, 2023). We describe the specific benchmarks and performances in the supplementary materials.
This work makes several key contributions. First, we introduce a novel distribution-based methodology for probing cultural alignment across languages, moving beyond direct survey approaches to better capture latent cultural values (Sorensen et al., 2024). Second, we provide the first systematic comparison of how improvements in scale and post-training affect cultural alignment and US-centric bias across English, Danish, Dutch, and Portuguese through a series of robust statistical models. Third, we release a dataset of model-generated responses across multiple languages and cultural contexts as well as our code, enabling future research into cultural alignment and bias. [^1] See github.com/jhrystrom/multicultural-alignment for code, data, and supplementary materials. Together, these contributions advance our understanding of how LLM development choices influence cultural representation while providing tools for ongoing investigation of these critical issues.
@@ 2 Measuring Cultural Alignment
Figure 2: Pearson correlations in value polarity scores across studied countries from the World Values Survey. Value polarity scores are the fraction of the population in favour of a given topic. All correlations are positive, with most being between 0.7–0.95.
This section defines ‘cultural alignment’ and how to measure it in LLMs. We conceptualise cultural alignment as reproducing distributions of values in a particular population. Then we show how to a) get a ground-truth distribution of values using the World Values Survey (§2.1) and b) elicit value distributions from LLMs (§2.2).
@@
Within a culture there will be a variety of stances to any particular topic. However, the distribution of stances will be characteristic among cultures. For instance, while around 8% of Danes are opposed to abortion, it is a much less contentious topic than in the US, where it’s close to 40% EVS/WVS (2022).
We posit that cultural alignment for a specific group of people can be operationalised as how well an LLM reproduces the distribution of values over a wide range of topics (Sorensen et al., 2024). Investigating distributions of responses differs from previous work that directly surveys the LLMs as regular participants (e.g., Cao et al., 2023). This approach also addresses concerns raised by Khan et al. (2025) about the instability of survey-based evaluations by focusing on aggregate distributions rather than individual responses and incorporating explicit controls for response consistency. Our goal is to get more naturalistic elicitations of the underlying values whilst avoiding sycophancy and response bias Sharma et al. (2023).
We operationalise reproduction as high correlations between value polarity scores: the fraction of people (or LLM responses) in favour of a topic in the population. Note, that we binarise issues to allow for simpler operationalisation. Below, we describe how we empirically estimate the value polarity score for the ground truth (§2.1) and LLMs (§2.2).
@@ 2.1 Ground Truth: World Values Survey
To get a ‘ground truth’ distribution of cultural values, we use the joint World Values Survey and European Values Survey (EVS; EVS/WVS, 2022). These surveys cover adults across 92 countries with samples that are nationally representative for gender, age, education, and religion. The surveys’ broad coverage enables cross-cultural comparability for the many countries covered by the surveys, though some scholars note challenges in ensuring response comparability across countries (Alemán and Woods, 2016). The WVS provides both country and language identifiers for each respondent, allowing us to define populations either as citizens of a country or speakers of a language using the same underlying respondent-level data.
We select questions with binary agree/disagree or rating scale formats that allow clear classification of positive vs. negative stances, excluding questions with multiple categorical response options (see the supplementary materials for the full list of questions). These questions span environment, work, family, politics, religion, and security. We convert responses to binary indicators by determining whether each response indicates support for the measured construct, with custom coding to handle the various question formats and reverse-scored items. Finally, we calculate the value polarity score as the demographically weighted proportion of respondents with affirmative stances. Formally, we can define the value polarity score for a given population, $\mathcal{P}$ (e.g., citizens in a country or speakers of a language) and topic, $q$, (i.e., question within the EVS/WVS) as shown in Eq. 1:
$\mathrm{VPS}_{\mathcal{P},q}=\sum_{i\in\mathcal{P}}\frac{w_{i}}{\sum_{j\in\mathcal{P}_{q}}w_{j}}\,A_{i,q}$ (1)
Here, $A_{i,q}$ is a binary indicator of whether participant $i$ has a positive stance on topic $q$, $w_{i}$ represents the survey-provided demographic weights, and $\mathcal{P}_{q}$ denotes respondents in population $\mathcal{P}$ who answered question $q$. The first term normalises the weights to account for missing responses and enables aggregation across any definition of a population (e.g., residents in a country, speakers of a language, etc.).
For example, if 80% of Danish respondents who answered the same-sex marriage question expressed support (after demographic reweighting), Denmark’s value polarity score for this topic would be 0.8. Thus, a culture’s values can be represented as a vector, where each element corresponds to a value polarity score for a specific topic.
@@ 2.2 Ecologically valid LLM responses
Developing ecologically valid prompts requires careful consideration. When evaluating LLM responses to value-laden topics, simply asking questions like “What proportion of people support topic X?” or “Do you support topic X?” proves inadequate (e.g., Rozado, 2024). Such direct approaches suffer from three key limitations: they generate false positives through excessive agreement, fail to reflect realistic usage patterns, and provide insufficient variation to assess cultural alignment (Röttger et al., 2024). They also struggle to capture instance-specific harms that emerge when systems misalign with users’ cultural contexts (Rauh et al., 2022).
Instead, we adopt an implicit approach by asking the model to generate responses from hypothetical respondents. For example, prompting “imagine surveying 10 random people on topic X. What are their responses?” This method reveals the model’s latent opinion distribution while avoiding the limitations of direct questioning. Details for prompt construction are provided in the supplementary materials.
@@
Having a method for eliciting distributions of values, the next step is to seed culture. One typical way of seeding a specific culture is to explicitly instruct the LLM either by mentioning a specific country (‘imagine surveying 10 random Americans’) or through describing specific personas (‘Imagine surveying a 85-year-old Danish woman…’; AlKhamissi et al., 2024). The problem with these demographic prompting approaches is that they stray from actual uses of LLMs. Users are unlikely to explicitly mention their demographic information or nationality (Zheng et al., 2023a).
Instead, we use language as a proxy for cultural origin. For instance, a prompt in Danish is assumed to come from a Dane. This approach creates an intentional distinction in our analysis: we can compare ‘language-level’ alignment (all speakers of a language globally) with ‘country-level’ alignment (all people from specific nations where that language is native). As argued by Havaldar et al. (2023), users speaking a particular language would expect culturally appropriate responses in that language. For languages spoken in multiple countries, this approach is intentionally ambiguous. The ambiguity allows us to elicit the underlying ‘default’ alignment rather than the general ability to emulate cultures (Tao et al., 2024). We validate this approach by showing that LLM responses exhibit significantly lower self-consistency between languages compared to within languages, demonstrating that language impacts output (see the supplementary materials). To create prompts across languages, we use gpt-3.5-turbo to translate our original English prompts. Although previous literature has shown strong translation capabilities in LLMs (Yan et al., 2024), we nonetheless manually verify the translations.
@@
Finally, to transform the LLMs’ hypothetical survey responses into vectors of stances, we use an LLM-as-a-judge approach Zheng et al. (2023b); Guerdan et al. (2025). Specifically, we use gpt-4.1-mini (OpenAI et al., 2025) to label each substatement as either ‘pro’, ‘con’, or ‘null’ given the context of the topic and a representative pro and con statement (generated with an LLM and validated by the authors). We then calculate the proportion of ‘pro’ versus ‘con’ responses as the LLM’s value polarity score for the given statement. For instance, a response with seven ‘pro’, one ‘con’, and two ‘null’ statement would yield a value polarity score of 0.875 ($\frac{7}{8}$). A complete, unabridged example can be found in the supplementary materials. Formally, we label each substatement from the full set of hypothetical statements, $G_{q,g}$, for topic $q$ and generation $g$ as $r$. Furthermore, we label the classifier as $\ell(r)$. We then formalise the value polarity score for a given instance of a generation for a topic ($\mathrm{VPS}^{\mathrm{LLM}}_{q,g}$) as shown in Eq. 2:
$\mathrm{VPS}^{\mathrm{LLM}}_{q,g}=\frac{\sum_{r\in G_{q,g}}[\,\ell(r)=\mathrm{pro}\,]}{\sum_{r\in G_{q,g}}[\,\ell(r)\in\{\mathrm{pro},\mathrm{con}\}\,]},$ (2)
These scores are then compared against the value polarity scores from the WVS. Specifically, we calculate the Spearman rank correlation to obtain a measure of similarity between the LLMs’ responses and the value distributions of a given population.
To validate the LLM-as-judge, we manually annotate 200 statements. We iteratively refine the prompts and the LLM used until we reach satisfactory performance. We find a 91% agreement and a mean absolute error for value polarity of 4.5% over the dataset, ensuring consistent statistics between LLM and human annotation (Guerdan et al., 2025).
@@ 3 Experimental Setup
To investigate whether improving the multilingual capabilities of LLMs improves cultural alignment, we set up an experiment using a carefully chosen set of models and languages. We examine two different kinds of model improvements: scaling and commercial product development. These cases provide complementary perspectives on the effects of multilingual capabilities on cultural alignment. Scaling is the most well-studied path to improving LLMs (Kaplan et al., 2020; Ganguli et al., 2022). Commercial product development, on the other hand, comprises both scale and innovation in post-training to accommodate different pressures from capabilities, cost, and preferences (Kirk et al., 2024a). For scaling, we use the instruction-tuned Gemma models Gemma et al. (2024) and OLMo-2 models OLMo et al. (2025), while for product development, we use OpenAI’s turbo-series models OpenAI (2022); OpenAI et al. (2024a, b). We provide details of these model families in §3.1. A breakdown of the computational cost is in the supplementary materials.
@@
For the languages, we compare English with Danish, Dutch, and Portuguese. This set allows us to test multiple assumptions about cultural alignment. English represents a widely used case: it is a global language with speakers across many countries represented in the WVS (see Fig. 2). This diversity allows us to assess whether LLMs align more strongly with US values or those of other English-speaking nations.
Danish and Dutch serve as controlled test cases since they are primarily used in a single country. If cultural alignment stems from pre-training data, models should show strong Danish/Dutch cultural alignment when using these languages, despite their small share of training data (Kreutzer et al., 2022). Alternatively, if alignment emerges from post-training processes—which are predominantly English-based (Blevins and Zettlemoyer, 2022)–responses in these languages should align more with US values.
@@
For each language-model pair, we collect 300 prompt-response pairs to power our statistical analysis sufficiently (see §3.2). After filtering out responses that either lacked the required hypothetical survey format or were in a language other than the prompt, we obtained between 111–299 valid responses per combination. We calculate the correlation in value polarity scores at three levels: country (e.g., US or Denmark), language (pooling all speakers of a given language), and global (weighted values from all WVS/EVS participants).